OpenStudy (anonymous):

Am studying polar coordinates and need some help. i know the following conversion rules : x=rcosΔ y=rsinΔ tan Δ = y/x r^2 = x^2+y^2. Now my problem is how to convert from polar cordinates to rectagular coordinates. For instance how to convert the following: a) Δ = Alt+227 /2 b) r = sec Δ Can you help me understand it by showing me the procedure i must follow. thanks

OpenStudy (anonymous):

sorry for my post. How do i use the mathematical symbols? How can i edit my post?

OpenStudy (anonymous):

\[a) \theta=\pi/3\]

OpenStudy (anonymous):

b)\[b) r=\sec \theta\]

OpenStudy (anonymous):

r=sec pi/3=2, so x =2 cos pi/3 =2*1/2=1.. similarly find the value of y

OpenStudy (anonymous):

How did you get the answers; I need the procedure you followed to reach the answer

OpenStudy (anonymous):

x=r cos (theta) ,and y= r sin(theta) and what is r??? its r=sec (theta), where theta =pi/3 so r= sec pi/3 =2 now put the values of theta and r and get the ans..